40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 8 20:50:49 2011, sonali, Update, Green Locking, Power of the AUX laser increased. Graph3.pngETMY_beam_powers.png
    Reply  Thu Jul 14 02:32:11 2011, sonali, Update, Green Locking, Power of the AUX laser increased. 
       Reply  Fri Jul 15 13:48:56 2011, sonali, Update, Green Locking, Power of the AUX laser increased. ETMY_after_fibre_coupling_labelled.pdf
Message ID: 4965     Entry time: Thu Jul 14 02:32:11 2011     In reply to: 4958     Reply to this: 4973
Author: sonali 
Type: Update 
Category: Green Locking 
Subject: Power of the AUX laser increased. 

Quote:

The power of the beam which is being coupled into the optical fibre is measured to be between 159 mW to 164 mW (The power meter was showing fluctuating readings).

The power out of the beam coming out of the fibre far-end at the PSL table is measured to be 72 mW. Here, I have attached a picture of the beam paths of the ETMY table with the beams labelled with their respective powers.

 For the phase locking or beat note measuring we only need ~1 mW. Its a bad idea to send so much power into the fiber because of SBS and safety. The power should be lowered until the output at the PSL is < 2 mW. In terms of SNR, there's no advantage to use such high powers.

ELOG V3.1.3-